OpenStudy (marsxtc):

If x is an acute angle and sin x = 12/13, thne cos 2x equals? is it 119/169?

OpenStudy (anonymous):

cos2x = 1 - 2sin^2 x 1 - 2sin^2 x = 1 - 2(12/13)^2 1 - 2(12/13)^2 = 1 - 288/169 1 - 288/169 = 169/169 - 288/169 -119/169. You just forgot the negative there.